OpenStudy (anonymous):

IF IT HAS BEEN DETERMINED THAT THE PROBABILITY OF AN EARTHQUAKE OCCURRING ON A CERTAIN DAY IN A CERTAIN AREA IS .05 WHAT ARE THE ODDS IN FAVOR OF AN EARTHQUAKE?

OpenStudy (michele_laino):

@dan815 please help

OpenStudy (dan815):

What are the answer choices

OpenStudy (dan815):

do you want to just cover into a:b form

OpenStudy (dan815):

convert*

OpenStudy (dan815):

5/100 = 5:95= 1:19

OpenStudy (michele_laino):

yes! I think so: the requested answer is given by the ratio: probability to have an earthquake/ probability not to have an earthquake=5/95

hartnn (hartnn):

odds in favor = number of favorable events/ number of unfavorable events here, favorable events = 0.05 unfavorable events = 1-0.05 odds in favor will be = 0.05/(1-0.05)

OpenStudy (anonymous):

SORRY GUYS HAD TO TAKE THE DOG FOR HIS MORNING WALK THOUGHT I HAD TIME

Miracrown (miracrown):

May the odds be ever in your favor

OpenStudy (anonymous):

I HATE DOING THESE FREAKING PROBLEMS, I PASSED TRIG AND CALC WITH FLYING COLORS AND I'M STRUGGLING WITH THIS CRAP

OpenStudy (usukidoll):

that's why I bailed out... instructor change at the last minute

Miracrown (miracrown):

Wait ... I see an apple O.O

OpenStudy (anonymous):

HUH? LOL

OpenStudy (usukidoll):

I was supposed to have a good instructor for Elementary Probability Theory, but bailed because the last minute replacement was junk, so I completed my Calculus studies instead. :)

OpenStudy (anonymous):

I'M CONFUSED ON HOW DAN ARRIVED AT HIS RATIOS...

OpenStudy (dan815):

A:B is defined to be A/(A+B)